OpenStudy (anonymous):

\[Given u=e ^{rcos \theta}.\cos(rsin \theta), v=e ^{rcos \theta}.\sin(rsin \theta) prove that du/dr=1/r.dv/d \theta and dv/dr=-1/rdu/d \theta\]

OpenStudy (anonymous):

Dipti Do you want the Calculus or Algebra ?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

Okay \[u = e^{rcosa}*cos(r sina)\] \[\frac{du}{dr} = cos(rsina)(e^{rcosa} (cosa - rsina\frac{da}{sr})) - e^{rcosa}(sin(rsina)(sina + rcosa\frac{da}{dr}))\] I hate your Problems

OpenStudy (anonymous):

\[v= e^{rcosa}*sin(rsina)\] \[\frac{dv}{da} = e^{rcosa}(cos(rsina(\frac{dr}{da}sina + cosa r)) + sin(rsina)(e^{rcosa}(\frac{dr}{da} cosa - sina r))\]

OpenStudy (anonymous):

There you go ..please check them .. I hate them your problems

OpenStudy (anonymous):

thanks, i know you hate this problems , but i want the help for answer